Grady manufactures medical supply kits. Projected production for the first months of production are (hint: no beginning inventory for October):

Oct production 14,000

Nov production 32,000

There are four materials per kit at a cost of $2.25 each. The company desires to have enough materials on hand at month end to supply 20% of next month's production. Payments for material purchases are made 50% in the month purchased and the rest in the next month.

Each kit uses 0.45 hours of direct labor at $15.00 per hour, payable in the month worked. Indirect costs are fixed per month at $118,000 and includes $20,000 of depreciation, and are payable in the month incurred.

What should they budget for cash payments for production costs in October?
Answer #1

Calculate October cash payment

First we calculate October purchase

October purchase = (14000*4)+(32000*4*20%)-(14000*4*20%) = 70400*2.25 = 158400

Cash payment

Purchase (158400*50%) 79200
Direct labor (14000*.45*15) 94500
Indirect cost 118000
Less: Depreciation -20000
Cash payment 271700